3-5 Oslo Company prepared the following contribution format income statement based on a sales volume of...

3-5

Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units):

Sales $ 80,000
Variable expenses 52,000
Contribution margin 28,000
Fixed expenses 21,840
Net operating income $ 6,160

If sales decline to 900 units, what would be the net operating income?
Answer #1

Sales price per unit=80,000/1000=$80 per unit

Variable cost per unit=52,000/1000=$52 per unit

Sales(80*900) 72,000
Variable expenses(52*900) 46800
Contribution margin 25200
Fixed expenses 21840
Net operating income $3360.

NOTE:Total fixed costs and variable cost per unit do not change with change in units